On Monday, farmer Tom collected three times as many eggs as farmer jack. On Tuesday farmer Tom sold 160 eggs. Farmer Jack then had twice as many eggs as farmer Tom. How many eggs did farmer Jack have?
Answer by Boreal(15235)

x=Tom
Monday: T has x ; Jack has (1/3)x
Tuesday: T has x-160
(1/3)x = (x=160) 2
(1/3)x =2x -320
(-5/3)x= -320
(5/3)x =320
Multiply by 3
5x=960
x=192
Tom collected 192 eggs; Jack had 64
he sold 160 and had 32 left, which is half of Jack's
Jack had 64 eggs